OpenStudy (anonymous):

Factor the expression below. 100w2x2 - 64y2z2 Which of the following binomials is a factor of the expression? A. 10w2x2 + 8y2z2 B. 50wx + 32yz C. 10wx + 8yz D. 50w2x2 + 32y2z2

OpenStudy (mathstudent55):

Notice that 100 is a perfect square. Also, w^2 is a perfect square. The same goes for x^2, 64, y^2, and z^2. Since both sides of the minus signs are all perfect squares, you have a difference of squares. Do you remember how to factor a difference of squares?

OpenStudy (anonymous):

i do not

OpenStudy (mathstudent55):

Here it is: \( a^2 - b^2 = (a + b)(a - b)\)

OpenStudy (mathstudent55):

First, you need to see what the left term is the square of. What is the square root of 100w^2x^2?

OpenStudy (mathstudent55):

Then, what is the square root of 64y^2z^2?

OpenStudy (mathstudent55):

The square rot of the first term is 10wx The square root of the second term is 8yz Following the pattern I showed you above, the factoring is: \(100w^2x^2 - 64y^2z^2 = (10wx + 8yz)(10wx - 8 yz) \)

OpenStudy (anonymous):

thank you so much

OpenStudy (mathstudent55):

You're welcome.